OpenStudy (anonymous):

Simplify the expression involving exponents as much as possible. No negative exponents, please! x ^-5 times (x^3)^4

OpenStudy (anonymous):

\[x^{-5}x^{3*4}=x^{12-5}=x^{7}\]

jimthompson5910 (jim_thompson5910):

\[\large (x^{-5})(x^3)^4\] \[\large (x^{-5})(x^{3*4})\] \[\large (x^{-5})(x^{12})\] \[\large x^{-5+12}\] \[\large x^{7}\] So \[\large (x^{-5})(x^3)^4=x^7\]

OpenStudy (anonymous):

what about a^3/a^6 times a^3/a.... would you treat it like a regular fraction?

OpenStudy (anonymous):

ok thx